# Strong geometry: knots

### v1

@@ 3.2. General setting
We shall follow similar arguments as those in the rank $3$ case, but this time we need to handle linear and affine issues at the same time. Let $\boldsymbol{\alpha}_{s_{1}}=\boldsymbol{\alpha}(I_{1}),\boldsymbol{\alpha}_{s_{2}}=\boldsymbol{\alpha}(I_{2})$ and $\boldsymbol{\alpha}_{s_{3}}=\boldsymbol{\alpha}(I_{3})$.
By Lemma 1, we have
(5) $\chi_{\Lambda}(I_{1},I_{2},I_{3},J)=\Delta(x_{i_{7}},x_{i_{8}},x_{i_{9}},\boldsymbol{\alpha}(\boldsymbol{\alpha}_{s_{1}},\boldsymbol{\alpha}_{s_{2}},\boldsymbol{\alpha}_{s_{3}}))=-\Delta(\boldsymbol{\alpha}(\boldsymbol{\alpha}_{s_{1}},\boldsymbol{\alpha}_{s_{2}},\boldsymbol{\alpha}_{s_{3}}),x_{i_{7}},x_{i_{8}},x_{i_{9}}).$
We shall show that
(6) $\begin{split}\boldsymbol{\alpha}(\boldsymbol{\alpha}_{s_{1}},\boldsymbol{\alpha}_{s_{2}},\boldsymbol{\alpha}_{s_{3}})&=-\chi_{\omega}(a_{1},a_{2},a_{3})\boldsymbol{\omega}\\ &=-\Delta\big{(}(x_{i_{1}}-\omega)\wedge(x_{i_{2}}-\omega),(x_{i_{3}}-\omega)\wedge(x_{i_{4}}-\omega),(x_{i_{5}}-\omega)\wedge(x_{i_{6}}-\omega)\big{)}\boldsymbol{\omega}.\\ \end{split}$
The desired equality follows by combining (5) and (6).
Since $\boldsymbol{\alpha}_{s_{1}},\boldsymbol{\alpha}_{s_{2}}$ and $\boldsymbol{\alpha}_{s_{3}}$ contain the common point $\boldsymbol{\omega}$ then $\boldsymbol{\alpha}(s_{1},s_{2},s_{3})$ must be either $\boldsymbol{\omega}$ or $-\boldsymbol{\omega}$ (or trivial). We have that
$\mathsf{sign}(\langle\boldsymbol{\alpha}(s_{1},s_{2},s_{3}),\boldsymbol{\omega}\rangle)=\Delta(\boldsymbol{\alpha}_{s_{1}},\boldsymbol{\alpha}_{s_{2}},\boldsymbol{\alpha}_{s_{3}},\boldsymbol{\omega}).$
We show that
(7) $\Delta(\boldsymbol{\alpha}_{s_{1}},\boldsymbol{\alpha}_{s_{2}},\boldsymbol{\alpha}_{s_{3}},\boldsymbol{\omega})=-\chi_{\omega}(a_{1},a_{2},a_{3}).$
If $\boldsymbol{\alpha}_{s_{k}}=\mathbf{0}$ for some $1\leq k\leq 3$ then both sides of (7) are equal to 00. We thus suppose that $\boldsymbol{\alpha}_{s_{k}}\neq 0$, $1\leq k\leq 3$. Notice that both sides of Equation (7) are invariant by translations. Therefore, without loss of generality, we assume that $\omega=0$, and hence, $\boldsymbol{\omega}=\mathbf{e}_{1}=(1,0,0,0)$. In this setting, the $\boldsymbol{\alpha}_{s_{k}}$ are orthogonal to $\mathbf{e}_{1}$, we may thus write $\boldsymbol{\alpha}_{s_{k}}=(0,\tilde{\boldsymbol{\alpha}}_{s_{k}})$. Moreover, by definition, in the corresponding linear space $\mathbb{R}^{4}$, $\boldsymbol{\alpha}_{s_{1}}$ is the vector orthogonal to
$\left\langle\mathbf{e}_{1},\begin{pmatrix}1\\ x_{i_{1}}\end{pmatrix},\begin{pmatrix}1\\ x_{i_{2}}\end{pmatrix}\right\rangle=\left\langle\mathbf{e}_{1},\begin{pmatrix}0\\ x_{i_{1}}\end{pmatrix},\begin{pmatrix}0\\ x_{i_{2}}\end{pmatrix}\right\rangle,$
that is, $\tilde{\boldsymbol{\alpha}}_{s_{1}}$ is orthogonal to $\langle x_{i_{1}},x_{i_{2}}\rangle$ and such that
$\det\begin{pmatrix}1&1&1&0\\ 0&x_{i_{1}}&x_{i_{2}}&\tilde{\boldsymbol{\alpha}}_{s_{1}}\\ \end{pmatrix}=\det\begin{pmatrix}1&0&0&0\\ 0&x_{i_{1}}&x_{i_{2}}&\tilde{\boldsymbol{\alpha}}_{s_{1}}\\ \end{pmatrix}=\det\begin{pmatrix}x_{i_{1}}&x_{i_{2}}&\tilde{\boldsymbol{\alpha}}_{s_{1}}\end{pmatrix}>0.$
Therefore, we may deduce that $\tilde{\boldsymbol{\alpha}}_{s_{1}}=x_{i_{1}}\wedge x_{i_{2}}\in\mathbb{R}^{3}/\mathbb{R}_{+}$.
By proceeding with similar arguments, it can be showed that $\tilde{\boldsymbol{\alpha}}_{s_{2}}=x_{i_{3}}\wedge x_{i_{4}}$ and $\tilde{\boldsymbol{\alpha}}_{s_{3}}=x_{i_{5}}\wedge x_{i_{6}}$, see Figure 9. Finally we have that
$\Delta(\boldsymbol{\alpha}_{s_{1}},\boldsymbol{\alpha}_{s_{2}},\boldsymbol{\alpha}_{s_{3}},\boldsymbol{\omega})=\Delta\begin{pmatrix}0&0&0&1\\ x_{i_{1}}\wedge x_{i_{2}}&x_{i_{3}}\wedge x_{i_{4}}&x_{i_{5}}\wedge x_{i_{6}}&0\end{pmatrix}=-\chi_{\omega}(a_{1},a_{2},a_{3}).$
@@ 4.1. Gauss code
A Gauss code of an oriented knot diagram with $n$ crossings is constructed as follows. Label all the crossings of the diagram from $1$ to $n$ (in an arbitrary manner). The Gauss code is derived by walking the knot, starting at point $P$ of the diagram (picked arbitrarily and other than a crossing). As we follow the diagram, we record the crossings we encounter, by writing down the labels preceded with an ‘O’ or ‘U’ to indicate whether the curve goes Over or Under strand, see Figure 10.
@@
4.2. Main result
Let $K$ be a knot in $\mathbb{R}^{3}$. We may associate to $K$ a sphere shadow which is just the radial projection $\pi_{y_{0}}(K)$. We may suppose that such a shadow is regular in the sense that it avoids cusps and tangency points (this can be obtained my making some suitable local modifications to $K$ without changing its type). Let $z$ be a point in $\pi_{y_{0}}(K)$, suppose that
$z=\pi_{y_{0}}(y_{1})=\cdots=\pi_{y_{0}}(y_{k}),\ k\geq 1$
where $\|y_{1}\|<\cdots<\|y_{k}\|$ with $y_{1}$ is the nearest to $y_{0}$.
We say that $z$ is a simple intersection if $k=2$ and a multiple intersection if $k\geq 3$. We may also avoid multiple intersections $z$ by moving pieces of the shadow around $z$ properly. To this end, it is enough to consider $k-2$ concentric cercles $C_{k},\dots,C_{2}$ of radius $\epsilon_{k-2}>\cdots>\epsilon_{2}$ all of center $z$ and move the piece of arc $i=k,\dots,2$ inside $C_{i}$ to its boundary. Notice that these modifications do not change the type of $K$ and they are done orderly from the furthest to the nearest point to $y_{0}$, see Figure 14. We notice that these local diagrams arising on this way only depends on the order in which the different strands met as we turn around the intersction point.
Figure 14. Sequence of modifications done around a multiple intersection.
Therefore, sphere shadows can be thought of as a 4-regular map (i.e., an embedding of a 4-regular planar graph into $\mathbb{S}^{2}$). An spherical diagram of $K$ is obtained from such a shadow by endowing the under/over information to each vertex, see Figure 15.
Figure 15. A spherical diagram of the Trefoil.
We may now prove Theorem 1.
Proof of Theorem 1. Let $X=\{x_{0},\dots,x_{n-1}\}$ be a set of points in $\mathbb{R}^{3}$ in general position. Let $K$ be the polygon knot formed by segments $[x_{0},x_{1}]\cup[x_{1},x_{2}]\cup\cdots\cup[x_{n-1},x_{0}]$. We show that $\mathsf{SGeom}(M(X))$ determines uniquely the type of $K$. Let us consider the radial projection $\pi_{x_{0}}$ emitting from $x_{0}$ of $K$. Notice that the semi-open intervals $(x_{0},x_{1}]$ and $(x_{0},x_{n-1}]$ are each mapped into a single point. Since $K$ is polygonal then $\pi_{K}$ may have tangency point that can be fixed as described above, see Figure 16.
Figure 16. Possible tangencies (white circle) appearing in the projection $\pi_{x_{0}}(K)$. (Left) Tangency of an arc and ends of two arcs (Right) tangency of two ends of arc.
The shadow $\pi_{x_{0}}(K)$ may also have multiple points $z$ that can be modified as described above. We can keep record of where (and the order) of the new intersections are done if the the order of the points $y_{i}$ (with respect to the distance to $y_{0}$) having $\pi_{x_{0}}(y_{i})=z$ as image is known.
This radial projection does not arise a spherical diagram but rather a spherical line diagram, that is, an embedding of a planar graph with all the vertices having degree 4 except for two special ones of degree one, see Figure 17.
Figure 17. A spherical line diagram of a line eight-figure knot.
We may associate a Gauss path diagram to a spherical line diagram in a similar manner as a Gauss diagram is associated to a knot diagram. Take an oriented segment, say $S$, with an initial and end points. Walk along $S$ marking it with the labels for the crossings in the order they are encountered. Now draw arcs between the marks on $S$ that have the same label. Orient each arc from the over crossing to the under crossing in the spherical line diagram. Mark each arc with $+$ or -- according to handedness’ rule, see Figure 18.
Figure 18. The Gauss path diagram arising from the spherical line diagram given in Figure 17.
Claim 1.
If two polygonal knots have the same Gauss path diagram then they are isotopic.
@@ 4.2. Main result
We will show that Gauss path diagram determines the spherical line diagram (up to a planar isotopy). This might be done by slightly adapting the classic procedure to show that Gauss diagrams determine knot diagrams given in [6]. However, here we propose a new different approach. We shall see that any face of the spherical line diagram can be determined by using the Gauss path diagram. We shall show that the set of faces (and whether two of them share an edge) in the spherical line diagram correspond to a set of valid travels in the Gauss path diagram.
Let us consider a walk $W$ on the arcs on the boundary of a face (say, anticlockwise) in the spherical line diagram. Notice that we can run through an arc either following its direction or in opposite direction, see Figure 19.
@@
We associate to $W$ a travel $T$ along the Gauss path diagram. Turns of $W$ at a crossing correspond to jumps in arrows (from head/tail to tail/head) and each arc in $W$ corresponds to a piece of the oriented segment $S$ (of the Gauss path diagram). Travel $T$ may follow or not the direction of $S$ (depending on whether or not $W$ follows the direction of the corresponding arc). Since $W$ goes around a face then this correspondence obeys certain rules according to the sign of the crossing and from which strand (over/under) the walk comes from and goes to. Theses rules are illustrated in Figure 20.
Figure 20. Local turnings of a walk around a crossing and the corresponding local travel in the Gauss path diagram (in bold).
The travel $T$ constructed this way is called a valid travel in the Gauss path diagram. In fact, if the labels North, South, East and West, as well as the signs of each arrow in a Gauss path diagram are fixed then, a valid travel $T$ verifies the rule :
@@
Therefore, $W$ will be necessarily a walk of a face in the spherical line diagram. Moreover, two valid travels sharing a same piece of $S$ (with the same direction) correspond to two faces sharing an edge.
Finally, we observe that a valid travel going through an end of the path it goes back with opposite direction orientation, see Figure 22.
@@
We notice that any planar isotopy of line diagrams can be extended to an isotopy of knots by fixing $x_{0}$ and keeping the segments $[x_{n-1},x_{0}]$ and $[x_{0},x_{1}]$ straight while doing the isotopy. ∎
By the above claim, we only need to show that the Gauss path diagram can be completely recovered from the strong geometry $\mathsf{SGeom}_{\mathsf{Aff}}(X)=(M_{\mathsf{Aff}}(X),M_{\mathsf{Wedge}}(\mathsf{in}(X)))$. To this end, we show that $\mathsf{SGeom}(X)_{\mathsf{Aff}}$ determines the following three issues.

### v2

@@ 3.2. General setting
The proof runs similarly as the one for Proposition 1. The point is that the left hand term is the relative position between the hyperplane represented by $\alpha(y_{1},\dots,y_{d-1})$ and the positive intersection $\alpha(v_{1},\dots,v_{d-1})$ of the hyperplanes represented by the $v_{i}=\alpha(\omega,x_{i,1},\dots,x_{i,d-2})$. The proof consists in showing that the latter is precisely
$(-1)^{d+1}\tilde{\Delta}\Big{(}\tilde{\alpha}(x_{1,1},\dots,x_{1,d-2}),\dots,\tilde{\alpha}(x_{d-1,1},\dots,x_{d-1,d-2})\Big{)}\omega.$
Since the hyperplanes represented by the $\alpha(\omega,x_{i,1},\dots,x_{i,d-2})$ all contain $\omega$, their positive intersection $\alpha(v_{1},\dots,v_{d-1})$ is some $\eta\omega$, with $\eta\in\{-1,+1\}$ (unless it is trivial, in which case the proof is straightforward). By definition,
$\Delta\Big{(}\alpha(\omega,x_{1,1},\dots,x_{1,d-2}),\dots,\alpha(\omega,x_{d-1,1},\dots,x_{d-1,d-2}),\eta\omega\Big{)}=1$
so we simply want to compute
$\eta=\Delta\Big{(}\alpha(\omega,x_{1,1},\dots,x_{1,d-2}),\dots,\alpha(\omega,x_{d-1,1},\dots,x_{d-1,d-2}),\omega\Big{)}.$
Note that $\alpha(\omega,x_{i,1},\dots,x_{i,d-2})=\tilde{\alpha}(p(x_{i,1}),\dots,p(x_{i,d-2}))$ where $p$ denotes the orthogonal projection on $\omega^{\perp}$ (because a vector orthogonal to $\omega$ is orthogonal to some $x$ if and only if it is orthogonal to $p(x)$, and because the choice of convention for the orientation of $\omega^{\perp}$ was made to have no "--" sign appear here).
Therefore, we have
$\eta$ $=\Delta\Big{(}\tilde{\alpha}(p(x_{1,1}),\dots,p(x_{1,d-2})),\dots,\tilde{\alpha}(p(x_{d-1,1}),\dots,p(x_{d-1,d-2})),\omega\Big{)}$
$=(-1)^{d+1}\Delta\Big{(}\omega,\tilde{\alpha}(p(x_{1,1}),\dots,p(x_{1,d-2})),\dots,\tilde{\alpha}(p(x_{d-1,1}),\dots,p(x_{d-1,d-2}))\Big{)}$
$=(-1)^{d+1}\tilde{\Delta}\Big{(}\tilde{\alpha}(x_{1,1},\dots,x_{1,d-2}),\dots,\tilde{\alpha}(x_{d-1,1},\dots,x_{d-1,d-2})\Big{)}$
by definition of $\tilde{\Delta}$, and this is what we wanted to show. ∎
Let us rephrase Proposition 2 in terms of abstract chirotope. We restrict ourselves to the rank 444 case and state it in terms of projection from affine space (handy for the knot’s application).
Proposition 3.
Let $n\geq 1$ be an integer and let $X=(\omega=x_{0},x_{1},\dots,x_{n})$ be a tuple of points in $\mathbb{R}^{3}$.
Let $i_{0}=0$, $i_{1},\dots,i_{9}\in[n]^{9}$ and let $I_{1}=(i_{0},i_{1},i_{2})$, $I_{2}=(i_{0},i_{3},i_{4})$, $I_{3}=(i_{0},i_{5},i_{6})$ and $J=(i_{7},i_{8},i_{9})$. Let $a_{1}=(i_{1},i_{2}),a_{2}=(i_{3},i_{4}),a_{3}=(i_{5},i_{6})$. Then,
$\chi_{\wedge}(I_{1},I_{2},I_{3},J)=-\chi_{\Lambda,\omega}(a_{1},a_{2},a_{3})\chi(i_{0},i_{7},i_{8},i_{9})$
where $\chi_{\Lambda},\chi$ and $\chi_{\wedge,\omega}$ are the chirotopes of the rank 4 matroids $M_{\mathsf{Wedge}}(\mathsf{in}(X))$ and $M_{\mathsf{Lin}}(\mathsf{in}(X))=M_{\mathsf{Aff}}(X)$ and the rank 3 matroid $M_{\mathsf{Wit}_{\omega}}(X)$ respectively.
Here, $M_{\mathsf{Wit}_{\omega}}(X)$ is the witness oriented matroid of the configuration of points in the sphere centered at $\omega$ obtained by radial projection of $X$.
Proof.
Since the norms of the vectors are not taken into account in a linear matroid and we can project them on the sphere, this is simply expressing Proposition 2 by using one of the equivalent definitions of wedge chirotope. ∎
@@
Corollary 3.
Let $X$ and $X^{\prime}$ be two $n$-tuples in $\mathbb{R}^{4}$. If $M_{\mathsf{Wedge}}(X)$ and $M_{\mathsf{Wedge}}(X^{\prime})$ are isomorphic then $M_{\mathsf{Lin}}(X)$ and $M_{\mathsf{Lin}}(X^{\prime})$ also are.
Proof.
According to Proposition 1, $\chi(a\wedge b,a\wedge c,b\wedge c)=\chi(a,b,c)^{2}$. By plugging this into the equality of Proposition 3, we obtain
$\chi\big{(}(\alpha(x,a,b),\alpha(x,a,c),\alpha(x,b,c),\alpha(a,b,c)\big{)}$ $=\chi_{x}(a,b,c)^{2}\chi(x,a,b,c)$
$=\chi(x,a,b,c)^{2}\chi(x,a,b,c)$
$=\chi(x,a,b,c).$
@@ 4.1. Gauss code
A Gauss code of an oriented knot diagram with $n$ crossings is constructed as follows. Label all the crossings of the diagram from 111 to $n$ (in an arbitrary manner). The Gauss code is derived by walking the knot, starting at point $P$ of the diagram (picked arbitrarily and other than a crossing). As we follow the diagram, we record the crossings we encounter, by writing down the labels preceded with an ‘O’ or ‘U’ to indicate whether the curve goes Over or Under strand, see Figure 10.
@@
4.2. Knotoids
Knotoids are variants of knots that were first introduced by Turaev in [20]. The idea is to consider diagrams of segments with endpoints. Contrarily to knots, they are only defined as equivalence classes of diagrams, since all embedding of the interval in $\mathbb{R}^{3}$ are ambiant isotopic. More precisely, a knotoid is an equivalence class of knotoid diagrams. A knotoid diagram is the image of an application $[0,1]\to\mathbb{R}^{2}$ that satisfies the usual properties of a knot diagram, that is, each point has at most two preimages, called crossings, and there is a finite number of such crossings; besides, crossings happen between two transversal strands. We further suppose that the endpoints do not lie on crossings. The equivalence relationship considered here is the one generated by planar isotopy and the usual Reidemeister moves ; endpoints are not allowed to cross strands.
Figure 14. The knotoid called line eight-figure.
Similarly as we do for knots, we may associate to a knotoid diagram its Gauss diagram. Take an oriented segment, say $S$, with an initial and end points. Walk along $S$ marking it with the labels for the crossings in the order they are encountered as walking through the knotoid from one end to the other. Now draw arcs between the marks on $S$ that have the same label. Orient each arc from the over crossing to the under crossing in the knotoid. Mark each arc with $+$ or -- according to handedness’ rule, see Figure 15.
Figure 15. The Gauss diagram arising from knotoid given in Figure 14.
Lemma 3.
A knotoid diagram is determined by its Gauss diagram (up to a planar isotopy).
@@ 4.2. Knotoids
A proof might follow by slightly adapting the classic procedure showing that Gauss diagrams determine knot diagrams given in [16]. Nevertheless, we propose a new approach. We will see that any face of the knotoid diagram can be determined by using the Gauss diagram. We shall show that the set of faces (and whether two of them share an edge) in the knotoid diagram corresponds to the set of valid travels in the Gauss diagram.
Let us consider a walk $W$ on the arcs on the boundary of a face (say, counterclockwise) in the knotoid diagram. Notice that we can run through an arc either following its direction or in opposite direction, see Figure 16.
@@
We associate to $W$ a travel $T$ along the Gauss diagram. Turns of $W$ at a crossing correspond to jumps in arrows (from head/tail to tail/head) and each arc in $W$ corresponds to a piece of the oriented segment $S$ (of the Gauss diagram). Travel $T$ may follow or not the direction of $S$ (depending on whether or not $W$ follows the direction of the corresponding arc). Since $W$ goes around a face then this correspondence obeys certain rules according to the sign of the crossing and from which strand (over/under) the walk comes from and goes to. Theses rules are illustrated in Figure 17.
Figure 17. Local turnings of a walk around a crossing and the corresponding local travel in the Gauss diagram (in bold).
The travel $T$ constructed this way is called a valid travel in the Gauss diagram. In fact, if the labels North, South, East and West, as well as the signs of each arrow in a Gauss diagram are fixed then, a valid travel $T$ verifies the rule :
@@
Therefore, $W$ will be necessarily a walk of a face in the knotoid. Moreover, two valid travels sharing a same piece of $S$ (with opposite direction) correspond to two faces sharing an edge.
Finally, we observe that a valid travel going through an end of the line goes back with opposite orientation, see Figure 19.
@@
4.3. Main result
Let $K$ be a knot in $\mathbb{R}^{3}$ (not passing through $y_{0}$). We may associate to $K$ a sphere shadow which is just the radial projection $\pi_{y_{0}}(K)$. We may suppose that such a shadow is regular in the sense that it avoids cusps and tangency points (this can be obtained my making some suitable local modifications to $K$ without changing its type). Let $z$ be a point in $\pi_{y_{0}}(K)$, suppose that
$z=\pi_{y_{0}}(y_{1})=\cdots=\pi_{y_{0}}(y_{k}),\ k\geq 1.$
We say that $z$ is a simple intersection point if $k=2$ and a multiple intersection point if $k\geq 3$. We may avoid multiple intersection points by moving pieces of the shadow around $z$ properly. As for standard knots, doing such a perturbation does not change the knot type, however, in our case we need slightly more. We have to make sure that the strong geometry captures the information determining the possible diagrams.
The set of possible perturbations is completely determined by the local type of the multiple intersection, that is, the order (say counterclockwise) in which the strands appear as we rotate around the intersection point (because this determines the local diagram up to planar isotopy). Suppose that the strands are oriented inducing a ‘right’ and a ‘left’ side of the strand. Hence, each strand $a_{i}$ partitions the other $a_{j}$, $j\neq i$ into two sets, those crossing $a_{i}$ left-to-right and those crossing it right-to-left.
Figure 20. (Left) Two consecutive strands (bold strand crosses both right-to-left) (Center) Two consecutive strands (bold strand crosses both left-to-right) (Right) Two nonconsecutive strands (bold strand crosses left-to-right one of them and right-to-left the other).
We have that two strands are consecutive if and only they partition the other strands in equal or opposite sets. Therefore, we simply need to know the sign of each crossing in order to know what all the possible local perturbation. This piece of information will be shown to be induced by the strong geometry.
If the reader is not comfortable with these local permutations, instead, a notion of knot diagram allowing multiple crossings could be defined. The reasoning to check that the strong geometry gives the information about the strands around such multiple crossings would be essentially the same as above.
Therefore, sphere shadows can be thought of as 4-regular maps (i.e., an embedding of a 4-regular planar graph into $\mathbb{S}^{2}$). A spherical diagram of $K$ is obtained from such a shadow by endowing the under/over information to each vertex, see Figure 21.
Figure 21. A spherical diagram of the Trefoil.
We may now prove Theorem 1.
Proof of Theorem 1. Let $X=(x_{0},\dots,x_{n-1})$ be tuple of points in $\mathbb{R}^{3}$ in general position. Let $K$ be the polygonal knot formed by segments $[x_{0},x_{1}]\cup[x_{1},x_{2}]\cup\cdots\cup[x_{n-1},x_{0}]$. We show that $\mathsf{SGeom}(X)$ determines uniquely the type of $K$. Let us consider the radial projection $\pi_{x_{0}}(K-\{x_{0}\})$ of $K\-\{x_{0}\}$ emitted from $x_{0}$.
Remark 2.
(a) $\pi_{x_{0}}$ maps each semi-open interval $(x_{0},x_{1}]$ and $(x_{0},x_{n-1}]$ into a single point. We thus have that the corresponding spherical diagram induced by $\pi_{x_{0}}(K)$ is a knotoid, denoted by $\mathcal{K}$.
(b) Since $K$ is polygonal with points in generic position then the shadow $\pi_{x_{0}}(K-\{x_{0}\})$ has no tangency point.
(c) $\pi_{x_{0}}(K)$ may have multiple intersection points. These can be fixed/modified as explained above.
We notice that any planar isotopy between knotoids can be extended to an isotopy of knots by fixing $x_{0}$ and keeping the segments $[x_{n-1},x_{0}]$ and $[x_{0},x_{1}]$ straight while doing the isotopy.
Therefore, by Lemma 3, it is enough to show that the Gauss diagram associated to knotoid $\mathcal{K}$ is completely determined by $\mathsf{SGeom}_{\mathsf{Aff}}(X)=(M_{\mathsf{Aff}}(X),M_{\mathsf{Wedge}}(\mathsf{in}(X)))$ (implying that $K$ can also be uniquely reconstructed from such strong geometry). To this end, we show that $\mathsf{SGeom}_{\mathsf{Aff}}(X)$ determines the following two points.